# DREAM: Scalable Red Teaming for Text-to-Image Generative Systems via Distribution Modeling

### v1

@@ 3.1 Threat Model
We consider the red team to be a benign (non-malicious) model owner aiming to proactively identify safety vulnerabilities in his/her own T2I generative system. Specifically, their goal is to find a set of diverse and effective prompts that can elicit unsafe or policy-violating outputs, in order to assess and improve safety before real-world use. Following previous works [21, 49, 51], we assume the red teamer (1) has full control over his/her T2I generative system, such as requesting it with arbitrary prompt and receive the resulting image (or an all-black image if blocked by filters), or access to the model’s parameters and gradients, and (2) can leverage auxiliary models (e.g., open-sourced LLMs) for assistance and has moderate computational resources to fine-tune these models.
@@ 3.2 Formulation of Red Teaming
Despite the growing importance of red teaming in evaluating the safety of T2I generative models, existing literature [21, 28, 20] largely lacks a formal formulation of what the red teaming task fundamentally entails. This absence has led to fragmented understanding and inconsistent objectives, which limits both theoretical analysis and the principled design of scalable red teaming methods.
To bridge this gap, we present a formal definition of red teaming in this section. Let $\mathcal{X}$ and $\mathcal{Y}$ be the prompt space and image space of the target T2I generative system, respectively, we can draw the following definition:
@@
Let $G:\mathcal{X}\rightarrow\mathcal{Y}$ be a T2I system mapping a text prompt $x\in\mathcal{X}=\mathcal{V}^{T}$ to an image $y\in\mathcal{Y}$, where $\mathcal{V}$ and $T$ represents the full vocabulary and the maximum prompt length of the system, respectively. Red teaming aims to find a prompt subset $\mathcal{A}\subseteq\mathcal{X}$ such that:
@@
where $\mathcal{O}:\mathcal{Y}\rightarrow\{0,1\}$ is a binary oracle classifier that outputs $1$ if the image is unsafe and 00 if the image is safe or the request is denied by the equipped safety filter.
Intuitively, this definition formulates red teaming as a combinatorial subset discovery problem, whose aim is to identify all prompts that can trigger the T2I model to output unsafe content from the full prompt set $\mathcal{V}^{T}$. Note that the oracle function $\mathcal{O}$ is fundamentally unobservable in practice, as determining whether an image is “unsafe” is often vague, influenced by context, culture, and subjective interpretation [9]. As a practical alternative, red teaming methods rely on a surrogate scoring function $S:\mathcal{Y}\rightarrow\mathbb{R}$, which approximates the oracle with an objective score (e.g., the confidence score of an NSFW image detector). A prompt is deemed unsafe if its surrogate score is large enough to exceed a threshold $\tau$, yielding the surrogate unsafe set $\mathcal{A}_{\tau}:=\{x\in\mathcal{X}\mid S(G(x))\geq\tau\}$. While the surrogate formulation makes the task operational, obtaining the exact solution of the unsafe set $\mathcal{A}_{\tau}$ remains computationally intractable, as the task essentially reduces to a combinatorial searching problem over the prompt space $\mathcal{X}$, whose size grows exponentially with the prompt length $T$, i.e., $|\mathcal{X}|=|\mathcal{V}|^{T}$. In such combinatorial settings, exhaustive enumeration is the only general procedure that can ensure complete accuracy [54], yet it requires evaluating the surrogate score $S(G(x))$ for every enumerated $x\in\mathcal{V}^{T}$, making it computationally infeasible even for modest values of $T$. As a result, exact discovery is impractical except in trivial cases.
Fortunately, previous works have shown that exact recovery of $\mathcal{A}_{\tau}$ is often unnecessary. For instance, unlearning a moderate number of diverse and representative unsafe prompts is often sufficient to invalidate a much broader class of similar unsafe prompts [10, 14, 55, 56]. Consequently, the practical goal of red teaming shifts from full enumeration to the discovery of a representative and diverse subset $\hat{\mathcal{A}}\subseteq\mathcal{A}_{\tau}$, which captures a wide range of unsafe prompts while remaining computationally tractable to obtain.
@@
1:Input: Seed distribution $\pi(x)$, number of prompts $N$, max steps $T$, scoring function $S(\cdot)$, T2I model $G(\cdot)$, threshold $\tau$, update operator $\textsc{Update}(\cdot)$
@@ 3.3 Limitations of Previous Works
With an understanding of the red teaming task formulation, we now take a closer look at existing methods [20, 53, 28, 21]. While prior works differ substantially in their technical implementation, we distilled them into a unified, generic prompt-level discrete optimization paradigm, formalized in Alg. 1. Under this view, a red teaming algorithm begins with a seed prompt sampled from a seed distribution $\pi(x)$, and iteratively applies an Update operator guided by a scoring function $S(G(x))$, where $G(x)$ denotes the image generated by the T2I model. This loop continues until a generation crosses a threshold or a step budget is reached, at which point the final prompt is collected and the process resets. Note that the choice of seed prompt distribution, the Update operator, as well as the scoring function are all method-dependent. Despite empirical progress in uncovering unsafe prompts, we identify that this core algorithmic structure introduces two fundamental limitations, making them less suitable for scalable red teaming.
@@
First, the Update operator essentially performs discrete optimization, which is inherently difficult due to the discontinuous and non-smooth nature of the ill-posed loss landscape of the discrete prompt space [57]. In fact, how to accurately obtain prompt-level gradient remains an open challenge in existing literature [58, 59]. As such, most existing methods [44, 45, 20, 60] resort to token-level gradient replacement, where each token is iteratively and greedily updated based on its locally estimated gradient with respect to $S(G(x_{i}^{(t)}))$. However, this limits the search space to local neighborhoods around the initial seed, making the optimization process highly sensitive to initialization [61]. While recent methods attempt to broaden the search space by prompting LLMs to generate sentence-level paraphrases [21, 28], these approaches are largely heuristic, lack convergence guarantees, and often result in unstable training dynamics in practice.
@@
Second, one can easily observe from Alg. 1 that current methods are essentially operating at the individual prompt level: each run starts from a fresh seed prompt $x_{i}^{(0)}$, performs a local search trajectory $\{x_{i}^{(1)},x_{i}^{(2)},\cdots,x_{i}^{(t)}\}$, outputs $x_{i}^{(t)}$, and then discards all intermediate states but the final output before restarting the next run. This stateless fashion is naturally sub-optimal, as the algorithm would not accumulate any knowledge about explored regions or learn from past failures. Therefore, it may revisit similar attempt trajectories, re-try strategies that are known ineffective in previous runs and converging to familiar local optima, especially if the seed prompts are semantically or syntactically similar [21]. This makes the algorithm inefficient and yield highly similar prompts with limited marginal utility. Moreover, this inefficiency is exacerbated by the inherently slow convergence of discrete optimization. For example, P4D [20] and UnlearnDiffAtk [60] require roughly 3,000 rounds of model invocation and gradient updates to optimize a single prompt, taking about 30 minutes per prompt on an NVIDIA RTX A100 GPU. These deficiencies make it very difficult to be scaled up for large-scale red teaming.
@@ 4.1 Distributional Red Teaming via Energy-Based Modeling
Formally, let $q^{*}(x)$ denote the true (but unknown) distribution over the target model’s problematic prompts, i.e., the probabilistic distribution from which samples $x\in\hat{\mathcal{{A}}}$ are drawn. Our goal is to learn a probabilistic distribution $p_{\theta}(x)$ parameterized by $\theta$ (e.g., an autoregressive language model $p_{\theta}(x)=\prod_{t=1}^{T}p_{\theta}(x_{t}\mid x_{<t})$), such that $p_{\theta}(x)$ approximates $q^{*}(x)$ as possible, which can be characterized by the following Kullback–Leibler divergence [62] objective:
@@
This formulation has several desirable properties. First, by modeling the distribution $p_{\theta}(x)$, our method naturally converts the prompt-level discrete optimization into continuous optimization over model parameters $\theta$. Second, since the parameters encode the distribution over prompts, each parameter update accumulates knowledge about which types of prompts are more or less likely to trigger unsafe outputs, thus promoting exploration efficiency during training. Furthermore, it is totally feasible to initialize $p_{\theta}(x)$ with a pre-trained language model. As a result, our method inherits strong priors from large-scale human language data, which enables the model to understand and explore nuanced expressions, innuendos, and cultural references, which are subtle signals that typically require human-like common sense or contextual awareness and are often inaccessible to previous token-level search methods. Finally, once training is complete, sampling from the learned distribution $p_{\theta^{*}}(x)$ is efficient. An arbitrary number of diverse prompts can be generated efficiently via forward passes, without requiring iterative search or gradient updates. This property makes our approach particularly suitable for red teaming, where a large-scale of unsafe prompts (e.g., thousands) are required for safety assessment and downstream safety-tuning.
@@
Despite these promising properties, the objective in Eq. (1) remains particularly challenging to optimize in practice. The core difficulty lies in that the ground-truth distribution $q^{*}(x)$ is fundamentally unknown and there exists no readily available dataset that is sufficiently representative of the full support of the target model’s problematic prompts. This makes direct optimization (e.g., through fine-tuning with MLE [63]) impossible. Fortunately, recent advances in implicit generative modeling [64] provide a viable pathway to tackle this challenge. Specifically, results from the theory of energy-based models [29, 64] suggest that even in the absence of explicit samples, the target distribution $q^{*}(x)$ can be implicitly characterized with a properly defined energy function $E(x)$, which is a real-valued function that assigns lower values to more likely (or desirable) samples, and higher values otherwise. Then, the unknown distribution $q^{*}(x)$ can be expressed as a Boltzmann distribution [65] $q^{*}(x)={\exp(-\beta\cdot E(x))}/{Z}$, where $Z$ is a constant that normalizes the distribution [29, 64] and $\beta>0$ is a hyperparameter. Then, by plugging it into Eq.(1), we have:
$\arg\min_{\theta}D_{\mathrm{KL}}(p_{\theta}\,\|\,q^{*})=\sum_{x\in\mathcal{X}}p(x)\log\left(\frac{p(x)}{q(x)}\right)=$ (2)
$\mathbb{E}_{x\sim p_{\theta}}\left[\log\frac{p_{\theta}(x)}{q^{*}(x)}\right]=\mathbb{E}_{x\sim p_{\theta}}\left[E(x)+\frac{1}{\beta}\cdot\log p_{\theta}(x)\right].$
@@
The derivation above reduces the otherwise intractable KL divergence to two surprisingly simple yet intuitive components: the first is to minimize the energy function $E(x)$, i.e., to steer $\theta$ toward regions in the prompt space where the energy is low and thus more desirable. The second objective acts as a regularizer that penalizes low-entropy distributions by minimizing the log-likelihood $\mathbb{E}_{x\sim p_{\theta}}[\log p_{\theta}(x)]$, thereby avoiding degenerate solutions where the model collapses to a narrow set of prompts.
@@ 4.2 Energy Function Design
Vision-level Harmfulness Energy. The primary goal of $E$ is to guide the model toward the target model’s vulnerable prompt distribution $\mathcal{A}_{\tau}$. However, directly assessing the harmfulness of a text prompt $x$ is difficult as the risk often emerges only after it is rendered into an image. Therefore, we take a vision-level approach by evaluating the output image $y=G(x)$ instead of the prompt itself. Specifically, we employ BLIP-2 [66], a pretrained vision-language model with strong generalization across diverse image-text domains, to compute a vision-level harmfulness energy as part of the energy function. It assesses how the generated image is semantically aligned with a predefined harmful concept, and assign lower energy to those prompts that algin better with the harmful concept.
Formally, given a generated image $y=G(x)$ and a pre-defined target description $c$ (like “an image containing nudity”), the textual description $c$ is first sent to a pretrained language encoder $\mathcal{T}_{\phi}$ to obtain the sentence-level semantic embedding $t=\mathcal{T}_{\phi}(c)$. Then, the image $y$ is passed through a vision encoder followed by a specialized transformer module known as the Q-Former [66]. This module employs a set of query embeddings to interact with the visual features via cross-attention and finally extracts a set of latent tokens $\mathcal{I}_{\psi}(y)=\{z_{1},\dots,z_{k}\}$, each representing a different fine-grained aspect of the image in the same vision-language embedding space. Then, we define the alignment score as:
@@
[…]
This formulation brings three key benefits. First, BLIP-2 provides better generalization even under distribution shifts, such as stylized or non-photorealistic images, making the alignment score more reliable across visual domains [66, 67]. Second, the approach enables flexible red teaming through natural language descriptions. One can easily shift the target by modifying the concept prompt, e.g., replacing $c$ with “an image depicting violence scenes” to target violent content. When a small set of reference images is available, prompt tuning techniques can also be used to further refine and control the targeted concept [68]. Third, the alignment score is continuous, allowing small improvement in prompt effectiveness to be reflected. This supports more stable optimization than discrete (e.g. binary) success signals.
Prompt-level Diversity Energy. While the combination of the alignment energy and entropy regularization in Eq. (3) and Eq. (4) is effective in steering the model toward $q^{*}(x)$, we find that the entropy term alone is often insufficient to ensure diverse generations. This is because the target distribution $q^{*}(x)$ may itself be biased, e.g., certain keywords like “nude” and their variant sentences might dominate the probability mass. As a result, semantically distinct prompts with lower probability under $q^{*}(x)$ may remain largely unvisited in limited sampling iterations. To address this, we introduce a diversity energy term that explicitly encourages broader coverage within a limited sample budget. Let $\mathcal{E}_{\xi}(x)\in\mathbb{R}^{d}$ denote the sentence embedding of prompt $x$, obtained from a frozen pre-trained encoder (e.g., a sentence transformer [69]). Then, we define the prompt-level diversity energy as the expected pairwise similarity among prompt embeddings sampled from the current model distribution $p_{\theta}(x)$:
@@
This would explicitly encourage semantic diversity among generated prompts in limited sampling iterations, thus promoting broader exploration and reducing redundancy.

### v2

@@ 3.1 Threat Model
We consider the red team to be a benign (non-malicious) model owner aiming to proactively identify safety vulnerabilities in their own T2I generative system. Specifically, their goal is to find a set of diverse and effective prompts that can elicit unsafe or policy-violating outputs, in order to assess and improve safety before real-world use. Specifically, we assume the red teamer (1) has full control over their T2I generative system. They may request it with an arbitrary prompt and receive the resulting image (or an all-black image if blocked by filters), or access the model’s parameters and gradients. Note that our DREAM only requires query access to the target model, making it applicable to both white-box and black-box systems. (2) can leverage auxiliary models (e.g., open-source LLMs) for assistance and has moderate computational resources to fine-tune these models.
@@ 3.2 Formulation of Red Teaming
Despite the growing importance of red teaming in evaluating the safety of T2I generative models, existing literature [li2024art, mehrabi2024flirt, chin2024p4d] largely lacks a formal formulation that captures the fundamental nature of the red teaming task.
To bridge this gap, we present a formal definition of red teaming in this section. Let $\mathcal{X}$ and $\mathcal{Y}$ be the prompt space and image space of the target T2I generative system, respectively. We can draw the following definition:
@@
Let $G:\mathcal{X}\rightarrow\mathcal{Y}$ be a T2I system that maps a text prompt $x\in\mathcal{X}=\mathcal{V}^{T}$ to an image $y\in\mathcal{Y}$, where $\mathcal{V}$ and $T$ represent the full vocabulary and the maximum prompt length of the system, respectively. Red teaming aims to find a prompt subset $\mathcal{A}\subseteq\mathcal{X}$ such that:
@@
where $\mathcal{O}:\mathcal{Y}\rightarrow\{0,1\}$ is a binary oracle classifier that outputs $1$ if the image is unsafe, and 0 if the image is safe or the request is blocked by the built-in safety filter.
Intuitively, this definition formulates red teaming as a combinatorial subset discovery problem that aims to identify all prompts within the full prompt set $\mathcal{V}^{T}$ that can trigger the T2I system to produce unsafe content. Note that the oracle function $\mathcal{O}$ is fundamentally unobservable in practice, as determining whether an image is “unsafe” is often vague, influenced by context, culture, and subjective interpretation [schramowski2023sld]. As a practical alternative, red teaming methods rely on a surrogate scoring function $S:\mathcal{Y}\rightarrow\mathbb{R}$, which approximates the oracle with an objective score (e.g., the confidence score of an NSFW image detector). A prompt is deemed unsafe if its surrogate score is large enough to exceed a threshold $\tau$, yielding the surrogate unsafe set $\mathcal{A}_{\tau}:=\{x\in\mathcal{X}\mid S(G(x))\geq\tau\}$. While the surrogate formulation makes the task operational, obtaining the exact solution of the unsafe set $\mathcal{A}_{\tau}$ remains computationally intractable, as the task essentially reduces to a combinatorial search problem over the prompt space $\mathcal{X}$, whose size grows exponentially with the prompt length $T$, i.e., $|\mathcal{X}|=|\mathcal{V}|^{T}$. In such combinatorial settings, exhaustive enumeration is the only general procedure that can ensure complete accuracy [hartmanis1982computers], yet it requires evaluating the surrogate score $S(G(x))$ for every enumerated $x\in\mathcal{V}^{T}$, making it computationally infeasible even for modest values of $T$. As a result, exact discovery is impractical except in trivial cases.
Fortunately, previous works have shown that exact recovery of $\mathcal{A}_{\tau}$ is often unnecessary. For instance, unlearning a moderate number of diverse and representative unsafe prompts is often sufficient to invalidate a much broader class of similar unsafe prompts [gandikota2023erasing, gandikota2024unified, ribeiro2018semantically, pei2017deepxplore]. Consequently, the practical goal of red teaming shifts from full enumeration to the discovery of a representative and diverse subset $\hat{\mathcal{A}}\subseteq\mathcal{A}_{\tau}$, which captures a wide range of unsafe prompts while remaining computationally tractable to obtain.
@@
1:Input: Seed distribution $\pi(x)$, number of prompts NN, max steps $T$, scoring function $S(\cdot)$, target T2I system $G(\cdot)$, threshold $\tau$, update operator $\textsc{Update}(\cdot)$
@@ 3.3 Limitations of Previous Works
With an understanding of the red teaming task formulation, we now take a closer look at existing methods [chin2024p4d, tsai2024ring, mehrabi2024flirt, li2024art]. While prior works differ substantially in their technical implementation, we distilled them into a unified, generic prompt-level discrete optimization paradigm, formalized in Alg. 1. Under this view, a red teaming algorithm begins with a seed prompt sampled from a seed distribution $\pi(x)$, and iteratively applies an Update operator guided by a scoring function $S(G(x))$, where $G(x)$ denotes the image generated by the T2I model. This loop continues until a generation crosses a threshold or a step budget is reached, at which point the final prompt is collected and the process resets. Note that the seed prompt distribution, the Update operator, and the scoring function are all method-dependent. Despite empirical progress in uncovering unsafe prompts, we identify that this core algorithmic structure introduces two fundamental limitations, making these methods less suitable for scalable red teaming.
@@
First, the Update operator essentially performs discrete optimization, which is inherently difficult due to the discontinuous and non-smooth nature of the ill-posed loss landscape of the discrete prompt space [bengio2013estimating]. In fact, how to accurately obtain prompt-level gradient remains an open challenge in existing literature [wen2023hard, wang2021adversarial]. As such, some existing methods [yang2024sneakyprompt, yang2024mma] resort to token-level gradient replacement, where each token is iteratively and greedily updated based on its locally estimated gradient with respect to $S(G(x_{i}^{(t)}))$. However, this limits the search space to local neighborhoods around the initial seed, making the optimization process highly sensitive to initialization [jiang2025jailbreaking]. While recent methods attempt to broaden the search space by prompting LLMs to generate sentence-level paraphrases [li2024art, mehrabi2024flirt], these approaches are largely heuristic, lack convergence guarantees, and often result in unstable training dynamics in practice.
@@
Second, one can easily observe from Alg. 1 that current methods are essentially operating at the individual prompt level: each run starts from a fresh seed prompt $x_{i}^{(0)}$, performs a local search trajectory $\{x_{i}^{(1)},x_{i}^{(2)},\cdots,x_{i}^{(t)}\}$, outputs $x_{i}^{(t)}$, and then discards all intermediate states but the final output before restarting the next run. This stateless fashion is naturally sub-optimal, as the algorithm would not accumulate any knowledge about explored regions or learn from past failures. Therefore, it may revisit similar attempted trajectories, re-try strategies that are known to be ineffective in previous runs, and converge to familiar local optima, especially if the seed prompts are semantically or syntactically similar [li2024art]. This makes the algorithm inefficient and results in highly similar prompts with limited marginal utility. Moreover, this inefficiency is exacerbated by the inherently slow convergence of discrete optimization. For example, P4D [chin2024p4d] requires roughly 3,000 rounds of model invocation and gradient updates to optimize a single red teaming prompt, taking about 30 minutes per prompt on an NVIDIA RTX A100 GPU. These deficiencies make it very difficult to be scaled up for large-scale red teaming.
@@ 4.1 Distributional Red Teaming via Energy-Based Modeling
Formally, let $q^{*}(x)$ denote the true (but unknown) distribution over the target model’s problematic prompts, i.e., the probabilistic distribution from which samples $x\in\hat{\mathcal{{A}}}$ are drawn. Our goal is to learn a probabilistic distribution $p_{\theta}(x)$ parameterized by $\theta$ (e.g., an autoregressive language model $p_{\theta}(x)=\prod_{t=1}^{T}p_{\theta}(x_{t}\mid x_{<t})$), such that $p_{\theta}(x)$ approximates $q^{*}(x)$ as close as possible. This objective can be characterized by the following Kullback–Leibler divergence [kullback1951information]:
@@
This formulation has several desirable properties. First, by modeling the distribution $p_{\theta}(x)$, our method naturally converts the prompt-level discrete optimization into continuous optimization over model parameters $\theta$. Second, since the parameters encode the distribution over prompts, each parameter update accumulates knowledge about which types of prompts are more or less likely to trigger unsafe outputs, thus promoting exploration efficiency during training. Furthermore, it is totally feasible to initialize $p_{\theta}(x)$ with a pretrained language model. As a result, our method inherits strong priors from large-scale human language data, which enables the model to understand and explore nuanced expressions, innuendos, and cultural references, which are subtle signals that typically require human-like common sense or contextual awareness and are often inaccessible to previous token-level search methods. Finally, once training is complete, sampling from the learned distribution $p_{\theta^{*}}(x)$ is efficient. An arbitrary number of diverse prompts can be generated efficiently via forward passes, without requiring iterative search or gradient updates. This property makes our approach particularly suitable for red teaming, where a large number of unsafe prompts (e.g., thousands) are required for safety assessment and downstream safety-tuning.
@@
Despite these promising properties, the objective in Eq. (1) remains particularly challenging to optimize in practice. The core difficulty lies in the fact that the ground-truth distribution $q^{*}(x)$ is fundamentally unknown and there exists no readily available dataset that is sufficiently representative of the full support of the target model’s problematic prompts. This makes direct optimization (e.g., through fine-tuning with MLE [fisher1922mathematical]) impossible. Fortunately, recent advances in implicit generative modeling [du2019implicit] provide a viable pathway to tackle this challenge. Specifically, results from the theory of energy-based models [lecun2006tutorial, du2019implicit] suggest that even in the absence of explicit samples, the target distribution $q^{*}(x)$ can be implicitly characterized with a properly defined energy function $E(x)$, which is a real-valued function that assigns lower values to more likely (or desirable) samples, and higher values otherwise. Then, the unknown distribution $q^{*}(x)$ can be expressed as a Boltzmann distribution [boltzmann1872weitere] $q^{*}(x)={\exp(-\beta\cdot E(x))}/{Z}$, where $Z$ is a constant that normalizes the distribution [lecun2006tutorial, du2019implicit] and $\beta>0$ is a hyperparameter. Then, by plugging it into Eq. (1), we have:
$\arg\min_{\theta}D_{\mathrm{KL}}\!\big(p_{\theta}\,\|\,q^{*}\big)=\arg\min_{\theta}\mathbb{E}_{x\sim p_{\theta}}\!\left[\log\frac{p_{\theta}(x)}{q^{*}(x)}\right]$ (2)
$=\arg\min_{\theta}\mathbb{E}_{x\sim p_{\theta}}\!\left[E(x)+\frac{1}{\beta}\log p_{\theta}(x)\right].$
@@
The derivation above reduces the otherwise intractable KL divergence to two simple yet intuitive components: the first is to minimize the expected energy $\mathbb{E}_{x\sim p_{\theta}}[E(x)]$, thereby shaping the distribution toward lower-energy (and thus more desirable) regions of the prompt space. The second objective acts as an entropy regularizer that penalizes low-entropy distributions by minimizing the log-likelihood $\mathbb{E}_{x\sim p_{\theta}}[\log p_{\theta}(x)]$, thus avoiding degenerate solutions where the model collapses to a narrow set of prompts.
@@ 4.2 Energy Function Design
Vision-level Harmfulness Energy. The primary goal of $E$ is to guide the learned distribution $p_{\theta}(x)$ toward the target model’s vulnerable prompt distribution $\mathcal{A}_{\tau}$. However, directly assessing the harmfulness of a text prompt $x$ is difficult as the risk often emerges only after it is rendered into an image. Therefore, we take a vision-level approach by evaluating the output image $y=G(x)$ instead of the prompt itself. Specifically, we employ BLIP-2 [li2023blip], a pretrained vision-language model with strong generalization across diverse image-text domains, to compute a vision-level harmfulness energy as part of the energy function. It assesses how the generated image is semantically aligned with a predefined harmful concept, and assigns lower energy to prompts that align better with the harmful concept.
Formally, given a generated image $y=G(x)$ and a predefined target description $c$ (like “an image containing nudity”), the textual description $c$ is first sent to a pretrained language encoder $\mathcal{T}_{\phi}$ to obtain the sentence-level semantic embedding $t=\mathcal{T}_{\phi}(c)$. Then, the image $y$ is passed through a vision encoder followed by a specialized transformer module known as the Q-Former [li2023blip]. This module employs a set of query embeddings to interact with the visual features via cross-attention and finally extracts a set of latent tokens $\mathcal{I}_{\psi}(y)=\{z_{1},\dots,z_{k}\}$, each representing a different fine-grained aspect of the image in the same vision-language embedding space. Then, we define the alignment score as:
@@
where $\langle\cdot,\cdot\rangle$ and $\|\cdot\|$ denote the inner product and the Euclidean norm, respectively. $E_{\text{align}}(x)$ measures the cosine similarity between the resulting image and the textual embedding, where higher similarity indicates stronger alignment with the harmful concept and thus lower energy.
This formulation brings three key benefits. First, BLIP-2 provides better generalization even under distribution shifts, such as stylized or non-photorealistic images, making the alignment score more reliable across visual domains [li2023blip, li2023blip2]. Second, the approach enables flexible red teaming through natural language descriptions. One can easily shift the target by modifying the concept prompt, e.g., replacing $c$ with “an image depicting violent scenes” to target violent content. When a small set of reference images is available, prompt tuning techniques can also be used to further refine and control the targeted concept [schramowski2022can]. Third, the alignment score is continuous, allowing small improvements in prompt effectiveness to be reflected. This supports more stable optimization than discrete (e.g., binary) success signals.
Prompt-level Diversity Energy. While jointly optimizing the harmfulness energy in Eq. (3) and the entropy term $\mathbb{E}_{x\sim p_{\theta}}[\log p_{\theta}(x)]$ in the main objective (Eq. (LABEL:eq:entropy)) effectively guides the model toward $q^{*}(x)$, we observe that relying solely on this formulation tends to produce a learned distribution with limited semantic diversity. This is possibly because the target distribution $q^{*}(x)$ may itself be biased, e.g., certain keywords like “nude” and their semantically similar variants might dominate the probability mass. Consequently, semantically distinct prompts with lower probability under $q^{*}(x)$ may remain largely unvisited in limited sampling iterations. To address this, we introduce a diversity energy term that explicitly encourages broader coverage within a limited sample budget. Let $\mathcal{E}_{\xi}(x)\in\mathbb{R}^{d}$ denote the sentence embedding of prompt $x$, obtained from a frozen pretrained encoder (e.g., a sentence transformer [sentence-transformers]). Then, we define the prompt-level diversity energy as the expected pairwise similarity among prompt embeddings sampled from the current model distribution $p_{\theta}(x)$:
@@
This would explicitly encourage semantic diversity among generated prompts within limited sampling iterations, thus promoting broader exploration and reducing redundancy.